# Randomized block-Krylov subspace methods for low-rank approximation of matrix functions

### v1

@@ Abstract
The randomized SVD is a method to compute an inexpensive, yet accurate, low-rank approximation of a matrix. The algorithm assumes access to the matrix through matrix-vector products (matvecs). Therefore, when we would like to apply the randomized SVD to a matrix function, $f(\bm{A})$, one needs to approximate matvecs with $f(\bm{A})$ using some other algorithm, which is typically treated as a black-box. Chen and Hallman (SIMAX 2023) argued that, in the common setting where matvecs with $f(\bm{A})$ are approximated using Krylov subspace methods (KSMs), more efficient low-rank approximation is possible if we open this black-box. They present an alternative approach that significantly outperforms the naive combination of KSMs with the randomized SVD, although the method lacked theoretical justification. In this work, we take a closer look at the method, and provide strong and intuitive error bounds that justify its excellent performance for low-rank approximation of matrix functions.
@@ 1 Introduction
Computing and approximating matrix functions is an important task in many application areas, such as differential equations [22, 24], network analysis [11, 13], and machine learning [9, 36]. For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$ with eigenvalue decomposition
@@
However, instead of computing $f(\bm{A})$ exactly, in many cases we only need an approximation to $f(\bm{A})$, such as a low-rank approximation. For example, recent work has shown that low-rank approximations can be used to accurately estimate the trace or diagonal entries of $f(\bm{A})$ [3, 10, 26, 28, 31]. Furthermore, when $f(\bm{A})$ admits an accurate low-rank approximation $\bm{B}$, one can accurately and cheaply approximate matrix-vector products (matvecs) with $f(\bm{A})$ using the approximation
@@
Popular algorithms to compute cheap, yet accurate, low-rank approximations of matrices include the randomized SVD, randomized subspace iteration, and randomized block Krylov iteration. For an overview and analysis of these algorithms, we turn readers to [21, 29, 40]. These methods can be referred to as matrix-free, since they do not require explicit access to the matrix; they only require implicit access via matvecs. Furthermore, the dominant computational cost of these algorithms is usually the number of matvecs with the matrix that we want to approximate.
@@
When the matrix we wish to approximate is a matrix-function, $f(\bm{A})$, we do not have access to exact matvecs. However, we can typically compute approximate matvecs through some black-box method, such as the (block) Lanczos method, or another Krylov subspace method [18, 20, 23, 4]. These methods are themselves matrix-free, only requiring matvecs with $\bm{A}$ to approximation matvecs with $f(\bm{A})$.
@@
To summarize, the standard approach for efficiently computing a low-rank approximation to a matrix function, $f(\bm{A})$, combines two black-box methods: an outer-loop like the randomized SVD that requires matvecs with $f(\bm{A})$, and an inner-loop that computes each of these matvecs using multiple matvecs with $\bm{A}$ A broad theme of this paper and related work is that computational benefits can often be obtained by opening up these black-boxes. I.e., we can get away with fewer total matvecs with $\bm{A}$ to obtain a low-rank approximation to $f(\bm{A})$ if we shed the simple inner/outerloop approach. Such an approach was taken in [6] and [32, 33], where it is shown that algorithms like the randomized SVD and Nyström approximation for low-rank approximation, and Hutch++ for trace estimation, can benefit from such introspection.
@@
Concretely, our work builds on a recent paper of Chen and Hallman [6], which develops so-called “Krylov-aware” variant of the Hutch++ and A-Hutch++ algorithms for trace estimation [28, 31] . The Hutch++ algorithm uses the randomized SVD to compute a low-rank approximation, which in turn is used reduce the variance of a stochastic estimator for the trace of $f(\bm{A})$ [8, 42, 43]. Chen and Hallman develop a more efficient method to compute this low-rank approximation for matrix functions. Numerical experiments suggest that their method is significantly more efficient than naively implementing the randomized SVD on $f(\bm{A})$ with approximate matvecs obtained in a black-box way from a Krylov subspace method. Our main contribution is to strengthen our theoretical understanding of the method from [6] by providing strong and intuitive error bounds that justify its excellent empirical performance.
@@ 1.1 Notation
For a matrix $\bm{B}$ we denote by $\|\bm{B}\|_{\textsf{{F}}}$ and $\|\bm{B}\|_{2}$ the Frobenius norm and spectral norm, respectively. For a sketching matrix $\bm{\Omega}\in\mathbb{R}^{n\times\ell}$, where $\ell\geq k$, we define
@@
For any matrix $\bm{B}\in\mathbb{R}^{m\times n}$, the submatrix consisting of rows $r_{1}$ through $r_{2}$ and columns $c_{1}$ through $c_{2}$ is denoted by $\bm{B}_{r_{1}:r_{2},c_{1}:c_{2}}$. We write $\bm{B}_{:,c_{1}:c_{2}}$ and $\bm{B}_{r_{1}:r_{2},:}$ to denote $\bm{B}_{1:m,c_{1}:c_{2}}$ and $\bm{B}_{r_{1}:r_{2},1:n}$ respectively. We write $\llbracket\bm{B}\rrbracket_{k}$ to denote the best rank $k$ approximation of $\bm{B}$ obtained by truncating the SVD. By $\mathbb{P}_{s}$ we denote the set of all polynomials whose degree is at most $s$. The Krylov subspace $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$ is defined as
$\mathcal{K}_{s}(\bm{A},\bm{\Omega}):=\operatorname{range}\left(\begin{bmatrix}\bm{\Omega}&\bm{A}\bm{\Omega}&\cdots&\bm{A}^{s-1}\bm{\Omega}\end{bmatrix}\right).$
@@ 2 Krylov-aware low-rank approximation
We now describe and motivate the algorithm that we will analyse. Inspired by [6], we will call it Krylov-aware low-rank approximation. We begin with outlining the block-Lanczos algorithm to approximate matvecs with matrix functions. Next, we present how one would naively implement the randomized SVD on $f(\bm{A})$ using the block-Lanczos method to approximate matvecs. Finally, we present the alternative Krylov-aware algorithm and why this method allows us to gain efficiencies.
@@ 2.1 The block-Lanczos algorithm
Given an $n\times\ell$ matrix $\bm{\Omega}$, the block-Lanczos algorithm (Algorithm 1) can be used to iteratively obtain an orthonormal (graded) basis for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$. In particular, using $s$ block-matrix vector products with $\bm{A}$, the algorithm produces a basis $\bm{Q}_{s}$ and a block-tridiagonal matrix $\bm{T}_{s}$
$\bm{Q}_{s}=\begin{bmatrix}\bm{V}_{0}&\cdots&\bm{V}_{s-1}\end{bmatrix},\quad\bm{T}_{s}=\bm{Q}_{s}^{\textsf{{T}}}\bm{A}\bm{Q}_{s}=\operatorname{tridiag}\left(\begin{array}[]{c}\begin{array}[]{cccc}\bm{R}_{1}^{\textsf{{T}}}&\cdots&\bm{R}_{s-1}^{\textsf{{T}}}\end{array}\\ \begin{array}[]{ccccc}\bm{M}_{1}&\cdots&\cdots&\bm{M}_{s}\end{array}\\ \begin{array}[]{cccc}\bm{R}_{1}&\cdots&\bm{R}_{s-1}\end{array}\end{array}\right),$ (3)
@@
6:     $\bm{Y}=\bm{Y}-\sum_{j=0}^{i-1}\bm{V}_{j}\bm{V}_{j}^{\textsf{{T}}}\bm{Y}$triangleright reorthogonalize (repeat as needed)
@@
The block Lanczos algorithm can be used to approximate matvecs and quadratic forms with $f(\bm{A})$ using the approximations
@@
If $f$ is a low-degree polynomial, then the approximations Eqs. 4 and 5 are exact.111(4) and (5) are written out under the assumption that $\bm{\Omega}$ has rank $\ell$. If $\operatorname{rank}(\bm{\Omega})=r<\ell$, then the index set $1:\ell$ should be replaced with $1:r$ in both (4) and (5).
Lemma 2.1 ([6, Lemma 2.1]).
@@
It follows that Eqs. 4 and 5 are good approximations if $f$ is well approximated by polynomials. [^2] We note that for block-size $\ell>1$, the Krylov subspace is not equivalent to $\cup\{\operatorname{range}(p(\bm{A})\bm{\Omega}):p\in\mathbb{P}_{s-1}\}$, and bounds based on best approximation may be pessimistic due to this fact. In fact, deriving stronger bounds is an active area of research; see e.g. [5, 14, 15, 16, 17, 25]. In particular, one can obtain bounds in terms of the best polynomial approximation to $f$ on $[\lambda_{\min},\lambda_{\max}]$ [38, 30, 2].
@@ 2.2 The randomized SVD for matrix functions
The randomized SVD [21] is a simple and efficient algorithm to compute low-rank approximations of matrices that admit accurate low-rank approximations. The basic idea behind the randomized SVD is that if $\bm{\Omega}$ is a standard Gaussian random matrix, i.e. the entries in $\bm{\Omega}$ are independent identically distributed $N(0,1)$ random variables, then $\operatorname{range}(\bm{B}\bm{\Omega})$ should be reasonable approximation to the range of $\bm{B}$’s top singular vectors. Hence, projecting $\bm{B}$ onto $\operatorname{range}(\bm{B}\bm{\Omega})$ should yield a good approximation to $\bm{B}$. Algorithm 2 implements the randomized SVD on a symmetric matrix $\bm{B}$. The algorithm returns either the rank $\ell\geq k$ approximation $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or the rank $k$ approximation $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$, depending on the needs of the user.
@@
input: Symmetric $\bm{B}\in\mathbb{R}^{n\times n}$. Rank $k$. Oversampling parameter $\ell-k$.
@@
Typically, the dominant cost of Algorithm 2 is that computation of matvecs $\bm{B}$. We require $2\ell$ such matvecs: $\ell$ in 2 with $\bm{\Omega}$ and $\ell$ in 4 with $\bm{W}$. When Algorithm 2 is applied to a matrix function $\bm{B}=f(\bm{A})$ these matvecs cannot be performed exactly, but need to be approximated using, for example, the block Lanczos method discussed in the previous section. Algorithm 3 implements the randomized SVD applied to $f(\bm{A})$ with approximate matvecs using the block Lanczos method. The cost is now $(s+r)\ell$ matvecs with $\bm{A}$, where $s$ and $r$ should be set sufficiently large so that the approximations Eqs. 4 and 5 are accurate.
@@
input: Symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Oversampling parameter $\ell-k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Accuracy parameters $s$ and $r$.

### v2

@@ Abstract
The randomized SVD is a method to compute an inexpensive, yet accurate, low-rank approximation of a matrix. The algorithm assumes access to the matrix through matrix-vector products (matvecs). Therefore, when we would like to apply the randomized SVD to a matrix function, $f(\bm{A})$, one needs to approximate matvecs with $f(\bm{A})$ using some other algorithm, which is typically treated as a black-box. Chen and Hallman (SIMAX 2023) argued that, in the common setting where matvecs with $f(\bm{A})$ are approximated using Krylov subspace methods (KSMs), a more efficient low-rank approximation is possible if we open this black-box. They present an alternative approach that significantly outperforms the naive combination of KSMs with the randomized SVD, although the method lacked theoretical justification. In this work, we take a closer look at the method, and provide strong and intuitive error bounds that justify its excellent performance for low-rank approximation of matrix functions.
Key words. Krylov subspace methods, Low-rank approximation, Randomized numerical linear algebra, Matrix functions
MSC Classes. 65F55, 65F60, 68W20
@@ 1 Introduction
Computing and approximating matrix functions is an important task in many application areas, such as differential equations [havel1994matrix, highamscaleandsquare], network analysis [estrada2000characterization, estradahigham], and machine learning [DongErikssonNickisch:2017, PleissJankowiakEriksson:2020]. For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$ with eigenvalue decomposition
@@
However, instead of computing $f(\bm{A})$ exactly, in many cases we only need an approximation to $f(\bm{A})$, such as a low-rank approximation. For example, recent work has shown that low-rank approximations can be used to accurately estimate the trace or diagonal entries of $f(\bm{A})$ [baston2022stochastic, epperly2023xtrace, jiang2021optimal, hpp, ahpp]. Furthermore, when $f(\bm{A})$ admits an accurate low-rank approximation $\bm{B}$, one can accurately and cheaply approximate matrix-vector products (matvecs) with $f(\bm{A})$ using the approximation
@@
Popular algorithms to compute cheap, yet accurate, low-rank approximations of matrices include the randomized SVD, randomized subspace iteration, and randomized block Krylov iteration. For an overview and analysis of these algorithms, we turn readers to [rsvd, MM15, tropp2023randomized]. These methods can be referred to as matrix-free, since they do not require explicit access to the matrix; they only require implicit access via matvecs. Furthermore, the dominant computational cost of these algorithms is usually the number of matvecs with the matrix that we want to approximate.
@@
When the matrix we wish to approximate is a matrix-function, $f(\bm{A})$, we do not have access to exact matvecs even in exact arithmetic. However, we can typically compute approximate matvecs through some black-box method, such as the (block) Lanczos method, or another Krylov subspace method [blocklanczos, guttel, functionsofmatrices, lanczosfunctionshandbook]. These methods are themselves matrix-free, only requiring matvecs with $\bm{A}$ to approximate matvecs with $f(\bm{A})$.
@@
To summarize, the standard approach for efficiently computing a low-rank approximation to a matrix function, $f(\bm{A})$, combines two black-box methods: an outer-loop like the randomized SVD that requires matvecs with $f(\bm{A})$, and an inner-loop that computes each of these matvecs using multiple matvecs with $\bm{A}$. A broad theme of this paper and related work is that computational benefits can often be obtained by opening up these black-boxes. I.e., we can get away with fewer total matvecs with $\bm{A}$ to obtain a low-rank approximation to $f(\bm{A})$ if we shed the simple inner/outerloop approach. Such an approach was taken in [chen_hallman_23] and [persson_kressner_23, funnystrom2], where it is shown that algorithms like the randomized SVD and Nyström approximation for low-rank approximation, and Hutch++ for trace estimation, can benefit from such introspection.
@@
Concretely, our work builds on a recent paper of Chen and Hallman [chen_hallman_23], which develops so-called “Krylov-aware” variant of the Hutch++ and A-Hutch++ algorithms for trace estimation [hpp, ahpp] . The Hutch++ algorithm uses the randomized SVD to compute a low-rank approximation, which in turn is used reduce the variance of a stochastic estimator for the trace of $f(\bm{A})$ [cortinoviskressner, ubarusaad, ubaru2017applications]. In these settings, even coarse approximations of $f(\bm{A})$ can yield computational benefits. This makes the task of computing such approximations relevant, even when $f(\bm{A})$ does not admit a highly accurate low-rank approximation. Chen and Hallman develop a more efficient method to compute this low-rank approximation for matrix functions and use the method in their version of the Hutch++ algorithm. Numerical experiments suggest that their method is significantly more efficient than naively implementing the randomized SVD on $f(\bm{A})$ with approximate matvecs obtained in a black-box way from a Krylov subspace method. Although their paper is framed around the task of estimating $\operatorname{tr}(f(\bm{A}))$, the key reason for the effectiveness of their method is that the low-rank approximation is more accurate than the randomized SVD. Consequently, to theoretically understand the approximation quality of their method, one must first understand the approximation quality of the low-rank approximation.
Our contributions are twofold. First, we strengthen the theoretical foundations of the low-rank approximation method introduced in [chen_hallman_23] by providing explicit and intuitive error bounds that justify its excellent empirical performance. To the best of our knowledge, this is the first work to provide low-rank approximation error bounds for general matrix functions; existing theory covers only operator monotone functions [persson_kressner_23, funnystrom2]. Classical theory for Krylov subspace methods have focused on (i) eigenpair approximation, (ii) approximations of the action $f(\bm{A})\bm{B}$, (iii) low-rank approximation of $\bm{A}$ itself. However, we are not aware of any results that translate existing results to low-rank approximation of matrix-functions. Moreover, our error bounds pave the way to obtain approximation guarantees for estimating $\operatorname{tr}(f(\bm{A}))$.
Second, our analysis builds on classical work on Krylov subspace methods for low-rank approximation [drineasipsenkontopoulouismail, rsvd, MM15, tropp2016randomized] and eigenvalue problems [shmuel, paige1971computation, saad1980rates, underwood1975iterative]. We aim to show how this rich literature can be used to obtain approximation guarantees for matrix functions. While we do not claim that the method analyzed here is optimal – in fact, in Section 6 we discuss a variant that performs even better in practice, but for which a precise error analysis is currently out of reach – we view this work as a first step toward a more general theory. Ultimately, we hope that future progress will make it possible to translate classical Krylov theory into general results for low-rank approximation of matrix functions, similar to how [funnystrom2] provided a general “black-box” theory that allowed classical results on low-rank approximation to be translated into approximation guarantees for operator monotone matrix functions.
Remark.
The bounds we prove in this work contain quantities that depend on the function $f$. In general these terms may be difficult to bound further without any additional assumptions on $f$. A key motivating example of the work by Chen and Hallman [chen_hallman_23] is the matrix exponential $\exp(\bm{A})$ for a symmetric matrix $\bm{A}$. Computing a low-rank approximation to $\exp(\bm{A})$ is a task that arises in numerous applications, ranging from network analysis [hpp], to quantum thermodynamics [chen_hallman_23, epperly2023xtrace], to solving PDEs [persson_kressner_23]. Note that the exponential has a decaying behavior. Hence, $\exp(\bm{A})$ may admit an accurate low-rank approximation, even when $\bm{A}$ does not. For these reasons, we provide explicit bounds for the exponential and focus our experiments on this case. [^1] Other functions, such as $\sqrt{x},\log(1+x),$ and $\frac{x}{x+1}$, are also relevant in this context. However, these functions are operator monotone, and in such cases, alternative methods are more suitable [persson_kressner_23, funnystrom2].
@@ 1.1 Notation
For a matrix $\bm{B}$ we denote by $\|\bm{B}\|_{\textsf{{F}}}$ and $\|\bm{B}\|_{2}$ the Frobenius norm and spectral norm, respectively. For a matrix $\bm{\Omega}\in\mathbb{R}^{n\times\ell}$ we define
@@
Hence, if $\ell\geq k$ and $\operatorname{rank}(\bm{\Omega}_{k})=k$ the pseudo-inverse $\bm{\Omega}_{k}^{\dagger}=\bm{\Omega}_{k}^{\textsf{{T}}}(\bm{\Omega}_{k}\bm{\Omega}_{k}^{\textsf{{T}}})^{-1}$ satisfies $\bm{\Omega}_{k}\bm{\Omega}_{k}^{\dagger}=\bm{I}$. For any matrix $\bm{B}\in\mathbb{R}^{m\times n}$, the submatrix consisting of rows $r_{1}$ through $r_{2}$ and columns $c_{1}$ through $c_{2}$ is denoted by $\bm{B}_{r_{1}:r_{2},c_{1}:c_{2}}$. We write $\bm{B}_{:,c_{1}:c_{2}}$ and $\bm{B}_{r_{1}:r_{2},:}$ to denote $\bm{B}_{1:m,c_{1}:c_{2}}$ and $\bm{B}_{r_{1}:r_{2},1:n}$ respectively. We write $\llbracket\bm{B}\rrbracket_{k}$ to denote the best rank (at most) $k$ approximation of $\bm{B}$ obtained by truncating the SVD. By $\mathbb{P}_{s}$ we denote the set of all scalar polynomials whose degree is at most $s$. The Krylov subspace $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$ is defined as
[…]
@@ 2 Krylov-aware low-rank approximation
We now describe and motivate the algorithm that we will analyse. Inspired by the work of Chen and Hallman [chen_hallman_23], we will call it Krylov-aware low-rank approximation. While the primary goal of Chen and Hallman’s work was trace estimation of matrix functions, a central component of their method is the Krylov-aware low-rank approximation method presented in this section.
We begin with outlining the block-Lanczos algorithm to approximate matvecs with matrix functions. Next, we present how one would naively implement the randomized SVD on $f(\bm{A})$ using the block-Lanczos method to approximate matvecs. Finally, we present the alternative Krylov-aware algorithm and why this method allows us to gain efficiencies.
@@ 2.1 The block-Lanczos algorithm
Given an $n\times\ell$ matrix $\bm{\Omega}$, the block-Lanczos algorithm (Algorithm 1) can be used to iteratively obtain an orthonormal (graded) basis for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$. In particular, using $s$ block-matvecs with $\bm{A}$, the algorithm produces an orthonormal basis $\bm{Q}_{s}$ and a block-tridiagonal matrix $\bm{T}_{s}$
$\bm{Q}_{s}=\begin{bmatrix}\bm{V}_{0}&\cdots&\bm{V}_{s-1}\end{bmatrix},\quad\bm{T}_{s}=\bm{Q}_{s}^{\textsf{{T}}}\bm{A}\bm{Q}_{s}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{bmatrix}\bm{M}_{1}&\bm{R}_{1}^{\textsf{{T}}}&&\\ \bm{R}_{1}&\ddots&\ddots&\\ &\ddots&\ddots&\bm{R}_{s-1}^{\textsf{{T}}}\\ &&\bm{R}_{s-1}&\bm{M}_{s}\end{bmatrix}}$, (3)
@@
6:  $\bm{Y}=\bm{Y}-\sum_{j=0}^{i-1}\bm{V}_{j}\bm{V}_{j}^{\textsf{{T}}}\bm{Y}$ $\triangleright$ reorthogonalize (repeat as needed)
@@
The Lanczos method has a rich history and has been studied extensively. Single-vector methods (that is, when $\ell=1$) have classically been studied in the context of approximating a small number of extreme eigenvalues [shmuel, paige1971computation, paige1972computational, scottthesis, krylovschur, wilkinson1971algebraic, thickrestart], and large-block methods have classically been studied in the context of low-rank approximation [drineasipsenkontopoulouismail, halkomartinssonshkolniskytygert, MM15, tropp2023randomized] and for approximating a few of the largest eigenvalues [cullum, saad1980rates, saad2011numerical, underwood1975iterative]. Moreover, the block-Lanczos algorithm can also be used to approximate matvecs and quadratic forms with $f(\bm{A})$ using the approximations [^2] Other approximations are also possible.
@@
Such approximations are often referred to as block-FOM for functions of matrices [OLeary1980, etna_vol47_pp100-126, frommer_lund_szyld_20] and block-Gauss-quadrature [golub_meurant_09, Zimmerling2025] respectively.
It is well-known that if $f$ is a low-degree polynomial, then the approximations Eqs. 4 and 5 are exact.333(4) and (5) are written out under the assumption that $\bm{\Omega}$ has rank $\ell$. If $\operatorname{rank}(\bm{\Omega})<\ell$, then the index set $1:\ell$ should be replaced with $1:\operatorname{rank}(\bm{\Omega})$ in both (4) and (5).
Lemma 2.1 (see e.g. [hochbruck_lubich, Section 2], [golub_meurant_09, Chapters 2-7], [chen_hallman_23, Section 3]).
@@
It follows that Eqs. 4 and 5 are good approximations if $f$ is well approximated by polynomials. [^4] We note that for block-size $\ell>1$, the Krylov subspace is not equivalent to $\cup\{\operatorname{range}(p(\bm{A})\bm{\Omega}):p\in\mathbb{P}_{s-1}\}$, and bounds based on best approximation may be pessimistic due to this fact. In fact, deriving stronger bounds is an active area of research; see e.g. [chen_greenbaum_musco_musco, frommer_schweitzer_guttel, frommer_lund_szyld_20, frommer_schweitzer, frommer_simoncini, hochbruck_lubich]. In particular, one can obtain bounds in terms of the best polynomial approximation to $f$ on $\lambda_{\min},\lambda_{\max}] [$Saad:1992, OrecchiaSachdevaVishnoi:2012, Amsel:2024].
@@ 2.2 The randomized SVD for matrix functions
The randomized SVD [rsvd] is a simple and efficient algorithm to compute low-rank approximations of matrices that admit accurate low-rank approximations. The basic idea behind the randomized SVD is that if $\bm{\Omega}$ is a standard Gaussian random matrix, i.e. the entries in $\bm{\Omega}$ are independent identically distributed $N(0,1)$ random variables, then $\operatorname{range}(\bm{B}\bm{\Omega})$ should be a reasonable approximation to the range of $\bm{B}$’s top singular vectors. Hence, projecting $\bm{B}$ onto $\operatorname{range}(\bm{B}\bm{\Omega})$ should yield a good approximation to $\bm{B}$. Algorithm 2 implements the randomized SVD on a symmetric matrix $\bm{B}$. If $\bm{\Omega}$ has $\ell\geq k$ columns, the algorithm returns either the rank $\ell\geq k$ approximation $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or the rank $k$ approximation $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$, depending on the needs of the user.
@@
input: Symmetric $\bm{B}\in\mathbb{R}^{n\times n}$. Rank $k$. Block-size $\ell\geq k$.
@@
Typically, the dominant cost of Algorithm 2 is the computation of matvecs with $\bm{B}$. We require $2\ell$ such matvecs: $\ell$ in 2 with $\bm{\Omega}$ and $\ell$ in 4 with $\bm{W}$. When Algorithm 2 is applied to a matrix function $\bm{B}=f(\bm{A})$ these matvecs cannot be performed exactly, but need to be approximated using, for example, the block-Lanczos method discussed in the previous section. Algorithm 3 implements the randomized SVD applied to $f(\bm{A})$ with approximate matvecs using the block-Lanczos method. The cost is now $(s+r)\ell$ matvecs with $\bm{A}$, where $s$ and $r$ should be set sufficiently large so that the approximations Eqs. 4 and 5 are accurate.
@@
input: Symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Block-size $\ell\geq k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Accuracy parameters $s$ and $r$.